OpenStudy (anonymous):

What's the answer to -14/9 - 7/45?

OpenStudy (anonymous):

You'll want to get a common denominator here. Your denominators are 9 and 45, so a denominator that is a multiple of both is 45. So leave the second fraction alone and multiply the top and bottom of the first fraction by 5: \[\frac{ -14 }{ 9 } - \frac{ 7 }{ 45 } = \frac{ -70 }{ 45 } - \frac{ 7 }{ 45 }\]

OpenStudy (anonymous):

-77/45 is the answer.

Directrix (directrix):

Violation --> Code of Conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution. @sunnygarg40